OpenStudy (anonymous):

How many solutions are there for the equation? x2 + 16 = 8x A. 2 B. 1 C. 0 How many solutions are there for the equation? x2 + 16 = 8x A. 2 B. 1 C. 0 @Mathematics

OpenStudy (anonymous):

AS DEGREE IS 2 SO ANSWERS WILL ALSO BE 2

OpenStudy (anonymous):

Not true. ^

OpenStudy (anonymous):

Y ESCOLAS?

hero (hero):

x^2-8x+16=0 x^2-4x-4x+16=0 x(x-4)-4(x-4)=0 (x-4)(x-4)=0 (x-4)^2=0 x-4 = 0 x=4 1 solution only

OpenStudy (anonymous):

Because your caps lock button is stuck :) But I'm assuming they're asking because there may be unreal solutions. Or if the parabola is tangent to the x intercept there will be 1 solution.

OpenStudy (anonymous):

IN THE ABOVE QUADRATIC EQUATION WE HAVE 2 ANSWERS. but in all cases it is not true the thing which u r saying because if u will solve it u will get two values of x